# OpenGVL - Benchmarking Visual Temporal Progress for Data Curation

### v2

@@ 1 Introduction
Advancements in hardware and modeling have accelerated progress in robotics. Various embodiments have recently been proposed with decreasing bill-of-material costs, leading to wider availability [1, 2, 3]. A variety of Vision-Language-Action (VLA) models are being created and open-sourced [4, 5, 6]. Furthermore, new benchmarks, repositories, and communities are being formed [7, 8]. These hardware innovations have led to different data collection approaches such as UMI and DexHub [9, 10]. However, this rapid progress is not matched by the availability of well-curated datasets. There are only a few large-scale datasets available, such as Agibot-World, OXE, and Droid [11, 12, 13]. Although these datasets are much larger than previously available ones, they remain an order of magnitude smaller than datasets used in vision or language domains [14, 15].
However, reduced entry barriers have led to wider adoption of different data collection methods and an increased propensity to share data. As of August 2025, more than $2.6$ million episodes were publicly shared on Hugging Face’s Dataset Hub alone. [^1] It is important to note that some datasets are copies of previously available datasets. This calls for building tools that allow efficient and cost-effective filtering of available data. Temporal prediction progress (general purpose reward functions) determines robots’ own proficiency at the specified task from their own observations [16, 17, 18]. Such ability can be repurposed to curate and filter already collected datasets [19].
Recently, Ma et al. [20] proposed a generative value function estimator (GVL) that leverages world knowledge embedded in VLMs to predict universal value functions and estimate task progress. To automatically measure episode or dataset quality, Ma et al. [20] introduced the Value-Order Correlation (VOC) metric, which exhibits useful characteristics for data curation applications.
Building on this foundation and motivated by the need for large-scale robotics datasets comparable to The Pile or C4 [14, 21], we develop an open-source temporal progress prediction system (OpenGVL) as a foundational tool for data management at scale. We replicate and extend the GVL approach for open-source models, creating the OpenGVL benchmark. Furthermore, we demonstrate how OpenGVL can serve as a practical tool for real-world data curation applications.
@@ 2 Related work
Learnable universal value functions and success detectors for robotics have been a long-standing challenge [22, 23, 17, 18, 24]. To formalize this concept, given a trajectory $\tau=(o_{1},\ldots,o_{T})$ with observations $o_{t}$, a value function $V:T\to\mathbb{R}$ assigns a scalar score to the entire trajectory, reflecting how well it achieves its goal. We define:
@@
where $r_{t}$ denotes the task success signal at step $t$ and $\gamma\in[0,1]$ is a discount factor. In practice, $V(\tau)$ serves as a temporal success measure that can be approximated from visual-language inputs. Initial works focused on sparse signals of task success and typically relied on training or fine-tuning base models for specific tasks or domains [24, 22]. Alakuijala et al. [25] and Zhang et al. [26] fine-tune a VLM with a sequential ranking objective to encourage later frames in the video to have higher rewards.
Recently, Ma et al. [20] proposed deriving fine-grained temporal success predictions through in-context learning. Leveraging advancements in VLMs, this approach naturally frames the problem as a trajectory, goal, and prediction setup where traditional training can be replaced with few-shot learning. In this setup, the VLM is provided with a few examples of trajectories along with their temporal value function progress and a trajectory to be evaluated. Ma et al. [20] showed that due to VLMs’ propensity for imitating behavioral patterns in context, shuffling input observations improves prediction quality.
@@ 3.1 Experimental Setup
Data Selection: We have targeted four initial datasets as a base for validation: nyu_door, berkeley_mvp, cmu_stretch, and nyu_franka [27, 11, 28]. These datasets represent diverse manipulation tasks spanning different robot embodiments with relatively low task complexity. From each dataset, we sampled $50$ episodes using the same episode indices. We consider two conditioning scenarios: zero-shot and two-shot, balancing open-source context capabilities with performance gains from additional episode examples [20]. Due to the limits on context length, for each episode we sample 15 random frames and shuffle both context and evaluation frames to provide equal context for all models. In the Appendix A, we share the full prompt used across all runs.
@@
Model Selection: We evaluate a comprehensive set of open-source VLMs spanning different parameter scales and architectural approaches. Our selection includes the Gemma-3 family [29], which provides models at 4B, 12B, and 27B parameter scales, and the Qwen2.5-VL-Instruct family [30] with 3B, 7B, and 32B parameter counts. Both families allow us to study the effect of model scaling on temporal progress prediction. Additionally, we include four models with integrated reasoning capabilities: GLM-4.1V-9B-Thinking [31] with 9B parameters, MiMo-VL-7B-RL-2508 [32] and Cosmos-Reason1-7B [33] with 7B parameters, and Kimi-VL-A3B [34] with 16B total parameters (3B active parameters), all of which incorporate thinking mechanisms that enable enhanced temporal reasoning through explicit reasoning steps. All selected models follow similar architectural paradigms with integrated vision and language encoders, enabling direct comparison of their temporal reasoning capabilities.
For comparison with proprietary models, we have also evaluated gpt-4o [35], gemini-2.5-flash-lite-preview-06-17, and gemini-2.5-pro [36] based on their context capabilities and previous performance. Since closed-source models are updated regularly, we initially tested their performance on unshuffled trajectories. We observed similar behavior to the versions evaluated in [20], confirming that these models tend to over-rely on temporal ordering cues in the provided context. Therefore, we evaluate all subsequent results using shuffled frames.
@@ 3.2 Benchmarking Open Source VLMs for GVL
Table 1 presents results for all models in the initial benchmark release under zero-shot and two-shot conditioning. The results show that the VLM scale improves temporal score quality. Both the largest Qwen and Gemma versions achieve similar scores with significant improvements over their smaller counterparts, while among reasoning models, MiMo-VL-7B-RL-2508 shows strong performance and GLM-4.1V-9B-Thinking demonstrates solid results, though Kimi-VL-A3B falls short despite good performance on other vision benchmarks [34, 31].
@@
Moreover, open-source counterparts reach only approximately $60–70\%$ of the performance of proprietary models’ upper bound scores [^2] The upper bound scores themselves are only a proxy to the (unobservable) ground truth.. This is a substantial gap compared to the smaller performance differences typically observed in text-only models. This finding demonstrates the importance of comprehensive VLM evaluation suites focused on robotics tasks [37] and highlights the much-needed progress in vision-language tasks requiring spatial reasoning.
@@ 4 Data curation in the wild
We show how OpenGVL can easily identify different dataset issues ranging from unclear task definitions and instructions to occluded sensors and failed/out-of-distribution examples that can disrupt training, as observed previously [20].
VLMs have already been employed to detect incorrect task instructions, one of the most pressing challenges since these often include ambiguous placeholders [6]. For example, SmolVLA sampled representative frames and provided them to the VLM alongside the original instructions. The VLM was prompted to produce a short, action-oriented sentence summarizing the behavior. We demonstrate that OpenGVL adds new dimensions to data curation by enabling effective filtering of problematic episodes or even entire datasets with ill-posed setups. We have identified three common issues with publicly shared data: (1) task definition problems, (2) labeling ambiguity, and (3) failed/out-of-distribution examples. In the following sections, we provide a detailed analysis of each category. We emphasize that we do not critique any specific submissions but rather highlight these challenges to improve future data collection efforts.

### v3

@@ 1 Introduction
Advancements in hardware and modeling have accelerated progress in robotics. Various embodiments have recently been proposed with decreasing bill-of-material costs, leading to wider availability [kocharm, so100, christoph2025orca]. A variety of Vision-Language-Action (VLA) models are being created and open-sourced [black2024pi, kim2024openvla, smolvla]. Furthermore, new benchmarks, repositories, and communities are being formed [cadene2024lerobot, ksim]. These hardware innovations have led to different data collection approaches such as UMI and DexHub [park2024dexhub, chi2024umi]. However, this rapid progress is not matched by the availability of well-curated datasets. There are only a few large-scale datasets available, such as Agibot-World, OXE, and Droid [openx, khazatsky2024droid, bu2025agibot]. Although these datasets are much larger than previously available ones, they remain an order of magnitude smaller than datasets used in vision or language domains [gao2020pile, raffel2019t5].
However, reduced entry barriers have led to wider adoption of different data collection methods and an increased propensity to share data. As of August 2025, more than $2.6$ million episodes were publicly shared on Hugging Face’s Dataset Hub alone. [^1] It is important to note that some datasets are copies of previously available datasets. This calls for building tools that allow efficient and cost-effective filtering of available data. Temporal prediction progress (general purpose reward functions) determines robots’ own proficiency at the specified task from their own observations [chen2021learning, guan2024tasksuccess, ma2022vip]. Such ability can be repurposed to curate and filter already collected datasets [cabi2020scalingdatadrivenroboticsreward].
Recently, ma2024vision proposed a generative value function estimator (GVL) that leverages world knowledge embedded in VLMs to predict universal value functions and estimate task progress. To automatically measure episode or dataset quality, ma2024vision introduced the Value-Order Correlation (VOC) metric, which exhibits useful characteristics for data curation applications.
Building on this foundation and motivated by the need for large-scale robotics datasets comparable to The Pile or C4 [gao2020pile, raffel2023t5], we develop an open-source temporal progress prediction system (OpenGVL) as a foundational tool for data management at scale. We replicate and extend the GVL approach for open-source models, creating the OpenGVL benchmark. Furthermore, we demonstrate how OpenGVL can serve as a practical tool for real-world data curation applications.
@@ 2 Related work
Learnable universal value functions and success detectors for robotics have been a long-standing challenge [du2023vision, rocamonde2023vision, guan2024tasksuccess, ma2022vip, yang2023rewardtune]. To formalize this concept, given a trajectory $\tau=(o_{1},\ldots,o_{T})$ with observations $o_{t}$, a value function $V:T\to\mathbb{R}$ assigns a scalar score to the entire trajectory, reflecting how well it achieves its goal. We define:
@@
where $r_{t}$ denotes the task success signal at step $t$ and $\gamma\in[0,1]$ is a discount factor. In practice, $V(\tau)$ serves as a temporal success measure that can be approximated from visual-language inputs. Initial works focused on sparse signals of task success and typically relied on training or fine-tuning base models for specific tasks or domains [yang2023rewardtune, du2023vision]. alakuijala2024vlc and zhang2025rewind fine-tune a VLM with a sequential ranking objective to encourage later frames in the video to have higher rewards.
Recently, ma2024vision proposed deriving fine-grained temporal success predictions through in-context learning. Leveraging advancements in VLMs, this approach naturally frames the problem as a trajectory, goal, and prediction setup where traditional training can be replaced with few-shot learning. In this setup, the VLM is provided with a few examples of trajectories along with their temporal value function progress and a trajectory to be evaluated. ma2024vision showed that due to VLMs’ propensity for imitating behavioral patterns in context, shuffling input observations improves prediction quality.
@@ 3.1 Experimental Setup
Data Selection: We have targeted four initial datasets as a base for validation: nyu_door, berkeley_mvp, cmu_stretch, and nyu_franka [Radosavovic2022, openx, bahl2023cmu]. These datasets represent diverse manipulation tasks spanning different robot embodiments with relatively low task complexity. From each dataset, we sampled $50$ episodes using the same episode indices. We consider two conditioning scenarios: zero-shot and two-shot, balancing open-source context capabilities with performance gains from additional episode examples [ma2024vision]. Due to the limits on context length, for each episode we sample 15 random frames and shuffle both context and evaluation frames to provide equal context for all models. In the Appendix A, we share the full prompt used across all runs.
@@
Model Selection: We evaluate a comprehensive set of open-source VLMs spanning different parameter scales and architectural approaches. Our selection includes the Gemma-3 family [team2025gemma], which provides models at 4B, 12B, and 27B parameter scales, and the Qwen2.5-VL-Instruct family [qwen2025qwen25] with 3B, 7B, and 32B parameter counts. Both families allow us to study the effect of model scaling on temporal progress prediction. Additionally, we include four models with integrated reasoning capabilities: GLM-4.1V-9B-Thinking [vteam2025glm] with 9B parameters, MiMo-VL-7B-RL-2508 [coreteam2025mimovltechnicalreport] and Cosmos-Reason1-7B [nvidia2025cosmos] with 7B parameters, and Kimi-VL-A3B [kimiteam2025kimivl] with 16B total parameters (3B active parameters), all of which incorporate thinking mechanisms that enable enhanced temporal reasoning through explicit reasoning steps. All selected models follow similar architectural paradigms with integrated vision and language encoders, enabling direct comparison of their temporal reasoning capabilities.
For comparison with proprietary models, we have also evaluated gpt-4o [achiam2023gpt], gemini-2.5-flash-lite-preview-06-17, and gemini-2.5-pro [gemini2025] based on their context capabilities and previous performance. Since closed-source models are updated regularly, we initially tested their performance on unshuffled trajectories. We observed similar behavior to the versions evaluated in [ma2024vision], confirming that these models tend to over-rely on temporal ordering cues in the provided context. Therefore, we evaluate all subsequent results using shuffled frames.
@@ 3.2 Benchmarking Open Source VLMs for GVL
Table 1 presents results for all models in the initial benchmark release under zero-shot and two-shot conditioning. The results show that the VLM scale improves temporal score quality. Both the largest Qwen and Gemma versions achieve similar scores with significant improvements over their smaller counterparts, while among reasoning models, MiMo-VL-7B-RL-2508 shows strong performance and GLM-4.1V-9B-Thinking demonstrates solid results, though Kimi-VL-A3B falls short despite good performance on other vision benchmarks [kimiteam2025kimivl, vteam2025glm].
@@
Moreover, open-source counterparts reach only approximately $60–70\%$ of the performance of proprietary models’ upper bound scores [^2] The upper bound scores themselves are only a proxy to the (unobservable) ground truth.. This is a substantial gap compared to the smaller performance differences typically observed in text-only models. This finding demonstrates the importance of comprehensive VLM evaluation suites focused on robotics tasks [team2025geminirob] and highlights the much-needed progress in vision-language tasks requiring spatial reasoning.
@@ 4 Data curation in the wild
We show how OpenGVL can easily identify different dataset issues ranging from unclear task definitions and instructions to occluded sensors and failed/out-of-distribution examples that can disrupt training, as observed previously [ma2024vision].
VLMs have already been employed to detect incorrect task instructions, one of the most pressing challenges since these often include ambiguous placeholders [smolvla]. For example, SmolVLA sampled representative frames and provided them to the VLM alongside the original instructions. The VLM was prompted to produce a short, action-oriented sentence summarizing the behavior. We demonstrate that OpenGVL adds new dimensions to data curation by enabling effective filtering of problematic episodes or even entire datasets with ill-posed setups. We have identified three common issues with publicly shared data: (1) task definition problems, (2) labeling ambiguity, and (3) failed/out-of-distribution examples. In the following sections, we provide a detailed analysis of each category. We emphasize that we do not critique any specific submissions but rather highlight these challenges to improve future data collection efforts.